Sodium hydroxide, NaOH, is dissolved in water to make up a solution that is 1.58×10−5M in NaOH. What is the pH of the solutionat 25.0∘C? Round the answer to three significant figures.
Answer #1

I You have 1.88810-SM NaOH poh = -logio Ców] = -logo C1:58X1D-5) por = 4.80 Also, we have the formula, PH+POH = 14 pH = 14-4.